# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ Appendix F Visualization
In this section, we provide visualization to explain our procedure and show our effectiveness directly. As shown in 7, 8, and 9, respectively. Figures exhibit the initial solution, optimal solution, and reported solutions (i.e., UDC-$\boldsymbol{x}_{2}$($\alpha$=50), UDC-$\boldsymbol{x}_{50}$($\alpha$=50), and UDC-$\boldsymbol{x}_{250}$($\alpha$=50)). The singular shape of initial solutions verifies our estimation that the dividing policy is mainly not to obtain a shorter predicted path, but to predict an initial solution that is easier to improve to the optimal solution by the following conquering stage. In addition, these examples also preliminarily verify that our trained UDC can handle large-scale TSP, CVRP, and OP problems well after several rounds of conquering.
@@ Appendix G Baselines & License
Classical solvers. For LKH3, We adopt the commonly used setting in the NCO methods [50] and implement the LKH3 in TSP, CVRP, OVRP, ATSP, and min-max mTSP based on the code from https://github.com/wouterkool/attention-learn-to-route. For min-max mTSP MAX_CANDIDATES is set to 6, and for OVRP VEHICLE is set to 100. In mim-max mTSP, the implemented HGA algorithm is given in https://github.com/Sasanm88/m-TSP. For OR-Tools, we implement the code in https://developers.google.com/optimization/pack/knapsack?hl=zh-cn for KP. EA4OP https://github.com/gkobeaga/op-solver is also implemented in OP.
SL-based sub-path solvers SL-based sub-path solvers contain the LEHD [8] and BQ [14]. In addressing OP by BQ, we use the reported gap to EA4OP in the BQ paper.
RL-based constructive solvers This paper involves AM [50], POMO [5], ELG [12], ICAM [16], MDAM [51], MatNet [70], Omni_VRP [75], and Sym-NCO [59] as baselines. Constructive solver DAN [45], DPN [60] and Equity-Transformer [67] are employed for min-max mTSP, and MatNet is specially employed for ATSP.
For POMO, ELG, Omni_VRP, and ICAM, We use multi-start with x8 augmentation for CO instances with size $N\leq 1,000$ and multi-start with no augmentation for instances with $N>1,000$ [16]. Specifically, the ELG aug×$\times$8 in Table 4 adopts a special setting given in its paper [76]. We adopt 200-augmentation for Sym-NCO on CO problems with size $N\leq 1,000$ and 100-augmentation on 2,000-node CO problems.
Heatmap-based For all involved heatmap-based solvers (i.e., Attn-MCTS [37], DIMES [35], DIFUSCO [39], and T2T [38]), we adopt the report results for comparison.
Neural divide-and-conquer methods We implement H-TSP [20] and GLOP [13] for test sets. For L2D, TAM, and SO, we list their reported results in comparisons. The random insertion in 16 is generated based on code in GLOP https://github.com/henry-yeh/GLOP
For datasets, we use the benchmark TSPLib, CVRPLib, and two provided test sets in DIMES https://github.com/DIMESTeam/DIMES?utm_source=catalyzex.com (MIS) and Omni_VRP https://github.com/RoyalSkye/Omni-VRP (TSP & CVRP).

### v2

@@ Appendix F Visualization
In this section, we provide visualization to explain our procedure and show our effectiveness directly. As shown in 7, 8, and 9, respectively. Figures exhibit the initial solution, optimal solution, and reported solutions (i.e., UDC-$\boldsymbol{x}_{2}$($\alpha$=50), UDC-$\boldsymbol{x}_{50}$($\alpha$=50), and UDC-$\boldsymbol{x}_{250}$($\alpha$=50)). The singular shape of initial solutions verifies that the dividing policy is mainly not to obtain a shorter predicted path, but to predict an initial solution that is easier to improve to the optimal solution by the following two conquering stages. In addition, these examples also preliminarily verify that our trained UDC can handle large-scale TSP, large-scale CVRP, and large-scale OP well after several rounds of conquering.
@@ Appendix G Baselines & License
Classical solvers. For LKH3, We adopt the commonly used setting in NCO methods [50] and implement the LKH3 in TSP, CVRP, OVRP, and min-max mTSP based on the code from https://github.com/wouterkool/attention-learn-to-route. For min-max mTSP, the parameter MAX_CANDIDATES is set to 6, and for OVRP VEHICLE is set to 100. In mim-max mTSP, the implemented HGA algorithm is given in https://github.com/Sasanm88/m-TSP. For OR-Tools, we implement the code in https://developers.google.com/optimization/pack/knapsack?hl=zh-cn for KP. EA4OP https://github.com/gkobeaga/op-solver is also implemented as an OP baseline and KaMIS https://github.com/KarlsruheMIS/KaMIS is implemented for MIS.
SL-based sub-path solvers SL-based sub-path solvers contain the LEHD [8] and BQ [14]. In addressing OP by BQ, we fail to get legal solutions so we use the reported gap to EA4OP.
RL-based constructive solvers This paper involves AM [50], POMO [5], ELG [12], ICAM [16], MDAM [51], MatNet [41], Omni_VRP [79], and Sym-NCO [61] as baselines. Constructive solver DAN [45], DPN [63] and Equity-Transformer [72] are employed for min-max mTSP, and MatNet is specially employed for ATSP. For POMO, ELG, Omni_VRP, and ICAM, We use the multi-start with x8 augmentation setting for CO instances with size $N\leq 1,000$ and multi-start with no augmentation for instances with $N>1,000$ [16]. Specifically, the ELG aug×$\times$8 in Table 4 adopts a special setting given in its paper [80]. We adopt 200-augmentation for Sym-NCO on CO problems with size $N\leq 1,000$ and 100-augmentation on 2,000-node CO problems.
Heatmap-based We fail to implement all involved heatmap-based solvers (i.e., Attn-MCTS [37], DIMES [35], DIFUSCO [39], and T2T [65]), so we adopt the report results for comparison.
Neural divide-and-conquer methods We implement H-TSP [20] and GLOP [13] as baselines. For L2D, TAM, and SO, we list their reported results in comparisons. The random insertion in 17 is generated based on code in GLOP https://github.com/henry-yeh/GLOP
Considering datasets, we use the benchmark TSPLib, CVRPLib, two SNAP graphs, and two provided test sets in DIMES https://github.com/DIMESTeam/DIMES?utm_source=catalyzex.com (MIS) and Omni_VRP https://github.com/RoyalSkye/Omni-VRP (TSP & CVRP).
@@
SNAP (ego-Facebook & feather-lastfm-social) Dataset Available online